### v3

# Can Many-Shot In-Context Learning Help Long-Context LLM Judges? See More, Judge Better!

Abstract
Leveraging Large Language Models (LLMs) as judges for judging the performance of LLMs has recently garnered attention. However, this type of approach is affected by the potential biases in LLMs, raising concerns about the reliability of the evaluation results. To mitigate this issue, we propose and study two versions of many-shot in-context prompts, which rely on two existing settings of many-shot ICL for helping GPT-4o-as-a-Judge in single answer grading to mitigate the potential biases in LLMs, Reinforced ICL and Unsupervised ICL. Concretely, the former utilizes in-context examples with model-generated rationales, and the latter without. Based on the designed prompts, we investigate the impact of scaling the number of in-context examples on the consistency and quality of the judgment results. Furthermore, we reveal the symbol bias hidden in the pairwise comparison of GPT-4o-as-a-Judge and propose a simple yet effective approach to mitigate it. Experimental results show that advanced long-context LLMs, such as GPT-4o, perform better in the many-shot regime than in the zero-shot regime. Meanwhile, the experimental results further verify the effectiveness of the symbol bias mitigation approach. The code and data are released in https://github.com/nick7nlp/SeeMoreJudgeBetter.
Can Many-Shot In-Context Learning Help Long-Context LLM Judges?
See More, Judge Better!
Mingyang Song, Mao Zheng, Xuan Luo Tencent Hunyuan nickmysong@tencent.com
1 Introduction
LLMs such as GPT-4 OpenAI (2023), Gemini 1.5 Reid et al. (2024), and Claude3 Anthropic (2024) have demonstrated remarkable capabilities across a wide range of Natural Language Processing (NLP) tasks, becoming integral tools in various applications. The rapid advancement of LLMs Chowdhery et al. (2023) underscores the critical need to evaluate their alignment with human intent in generated responses. Therefore, evaluation has emerged as a crucial research area pivotal to the success of LLMs Chang et al. (2023).
LLMs like GPT-4 have shown exceptional performance across various tasks, leading to their wide adoption as both evaluators Wang et al. (2023a); Fu et al. (2023a); Wang et al. (2023c); Zheng et al. (2023); Wang et al. (2023b); Chen et al. (2024) and annotators Peng et al. (2023). However, the reliability of LLMs as evaluators remains uncertain, given their sensitivity to textual instructions Xu et al. (2023); Turpin et al. and potential judgment biases Wang et al. (2023b); Zheng et al. (2023); Chen et al. (2024). To this end, researchers, such as Wang et al. (2023b), focus on proposing various methods to solve the potential biases that exist when LLMs act as judges, such as the positional bias.
Figure 1: Consistency between different versions of judgment results by adopting GPT-4o as a zero-shot judge. $\hat{v}_{1}$ and $\hat{v}_{2}$ are the results based on Prompt(A) in Table 3. $v_{1}$, $v_{2}$, and $v_{3}$ are results based on Prompt(B) in Table 3. Prompts (A) and (B) differ in whether to output the rating first or later. The consistency evaluations show that Prompt (A) and (B) almost obtain the agreement results, but the latter is convenient for constructing many-shot in-context examples, so we adopt the latter generated rationales in this study. $v_{1}$ vs. $v_{2}$ denotes comparing the first and second versions of evaluations. $v_{1}$ vs. $v_{2}$ vs. $v_{3}$ denotes the comparison between the three versions of evaluations.
When human and LLM judges grade answers involving model-generated chain-of-thought rationales, consistently and precisely assigning the same score in 1-10 across multiple judgments is challenging, even if a reference correct answer is provided and each score in 1-10 has a particular and comprehensive description. Meanwhile, it is impossible to cover all situations in the descriptions of scores. Nonetheless, existing LLM-as-a-Judge approaches only adopt prompts with several aspects of constraints, which is undoubtedly very difficult because LLMs do not know exactly what kind of answer corresponds to each score.
Newly expanded context windows of LLMs allow researchers to investigate ICL with more shots than the zero-shot and few-shot regimes, namely many-shot ICL. To fully investigate the many-shot ICL, Agarwal et al. (2024) propose two settings of many-shot ICL (i.e., Reinforced ICL and Unsupervised ICL) to explore scaling in-context learning to hundreds or thousands of in-context examples and find performance improvements across multiple tasks. More importantly, they show that using the many-shot in-context examples with chain-of-thought rationales generated through the zero-shot regime is effective, and the many-shot ICL may overcome the pre-training biases of LLMs, whereas few-shot ICL struggles. Therefore, the intuitive idea is to use the many-shot ICL, allowing LLM-as-a-Judge to see the zero-shot evaluations of similar questions and answers first and then scoring examples before scoring. Therefore, an interesting question arises:
  • Can many-shot in-context learning help long-context LLM judges, such as GPT-4o?
Motivated by prior findings and the above issue, we verify the consistency of the widely used prompts of LLM-as-a-Judge Zheng et al. (2023), as shown in Table 3. Concretely, the consistency experiments are based on the entire test set of GSM8K Cobbe et al. (2021), where the inference answers are obtained based on LLaMA3-70B (The pipeline as illustrated as in Figure 2 and details in §$\S$ 3). Figure 1 presents that the consistency between the two versions of evaluations is low, with nearly half of the ratings being inconsistent.
Inspired by the above phenomena and issues, in this paper, we investigate whether many-shot in-context learning helps long-context LLM judges. Motivated by Agarwal et al. (2024), we introduce two versions of many-shot ICL prompts via Reinforced ICL and Unsupervised ICL for LLM-as-a-Judge. The former utilizes in-context examples with model-generated rationales, and the latter removes rationales used in the former, details in §$\S$ A.1. Meanwhile, we reveal a novel symbol bias in GPT-4o-as-a-Judge and explore a simple approach for mitigating this issue. Experiments show that many-shot ICL can help GPT-4o-as-a-Jduge obtain higher quality and consistency evaluation results. As the number of in-context examples increases, the quality and consistency of evaluation improves significantly. Furthermore, we further verify the effectiveness of the proposed simple yet effective approach for mitigating the symbol bias in pairwise comparison of GPT-4o-as-a-Jduge.
In summary, our main contributions can be summarized as the following three-fold:
  • Firstly, we propose investigating many-shot ICL to help long-context LLM judges, such as GPT-4o. To the best of our knowledge, we are the first to explore and introduce many-shot in-context learning to assist long-context LLMs as judges in single answer grading of GPT-4o-as-a-Judge and scale the number of in-context examples to verify the effectiveness.
  • Second, inspired by the Counting-Stars benchmark Song et al. (2024), we reveal a novel potential bias in pairwise comparison of GPT-4o-as-a-Judge, namely symbol bias, and propose a simple yet effective approach to mitigate it.
  • Finally, through experimental results, many-shot ICL can help GPT-4o-as-a-Jduge obtain higher quality and consistency evaluation results. Meanwhile, we demonstrate the effectiveness of the symbol bias mitigation method.
Figure 2: The pipeline of the experiments.
Figure 3: Consistency between two versions of judgment results. Concretely, the bar corresponding to "0" on the x-axis represents the number of samples with consistent and inconsistent ratings in comparing evaluation results obtained twice using GPT-4o as the judge in the zero-shot regime. In addition, the zero-shot generated rationales are used for Reinforced ICL and appended to the prompt for Unsupervised ICL. The bar corresponding to "$2^{n}$" on the x-axis represents the consistency of using the GPT-4o as a judge in the many-shot Reinforced ICL regime.
2 GPT-4o as A Long-Context LLM Judge
In this section, we briefly describe the background of many-shot ICL, recall judgment biases, and introduce two variants of designed prompts for GPT-4o-as-a-Judge in this study.
2.1 Background of Many-Shot ICL
LLMs excel at few-shot in-context learning, which involves learning from a few input-output demonstrations (“shots”) provided in context at inference without weight updates Brown et al. (2020). Newly expanded long-context LLMs allow us to investigate ICL with hundreds of in-context examples Li et al. (2023); Agarwal et al. (2024). Due to some limitations, we only examine GPT-4o-as-a-Judge in a many-shot regime in this paper.
2.2 Recalling Judgment Biases
LLM judges possess potential biases, which have been widely explored Wang et al. (2023b); Wu and Aji (2023); Zheng et al. (2023); Chen et al. (2024). For example, positional bias refers to the phenomenon where, during pairwise comparisons, LLM judges tend to favor one side of a pair regardless of the actual quality of the answers. In this paper, we investigate leveraging many-shot ICL to help GPT-4o as a better Judge, maybe through reducing judgment biases.
2.3 Two GPT-4o-as-a-Judge Variations
Inspired by Zheng et al. (2023), we introduce two GPT-4o-as-a-Judge variations: single answer grading and pairwise comparison.
2.3.1 Single Answer Grading
In this variation, the GPT-4o Judge is required to assign a score with a reason to a single answer directly. Inspired by the recent work Agarwal et al. (2024), we adopt the Reinforced ICL, which uses model-generated rationales as the in-context examples. Meanwhile, we also design the prompt within the Unsupervised ICL. The 4-shot Reinforced ICL and 1-shot Unsupervised ICL append 4-shots with prompts are presented in Table 1 and Table 5.
2.3.2 Pairwise Comparison
The GPT-4o Judge is presented with a question and two answers and tasked with determining which is better or declaring a tie. The prompt used for this scenario is given in Table 3. In this work, we adopt the pairwise comparison of GPT-4o-as-a-Judge to evaluate the quality of the GPT-4o judgments results in the zero-shot regime.
Figure 4: Consistency between two versions of judgment results.
Figure 5: Consistency between two versions of judgment results.
Figure 6: Consistency between two versions of judgment results.
3 Experiments
In this section, we introduce the experimental settings, consistency evaluation in single answer grading of GPT-4o-as-a-Jduge, quality evaluation via pairwise comparison of GPT-4o-as-a-Jduge, and reveal the symbol bias in GPT-4o-as-a-Jduge.
3.1 Experimental Settings
To analyze the GPT-4o judge in the many-shot regime, we use LLaMA3-70B to generate answers for each question in GSM8K Cobbe et al. (2021) with a temperature of 0.7. Details of the used benchmark are provided in §$\S$ A.3 (Appendix). There are two reasons for using LLaMA3 to infer the answers to the problems in GSM8K. First, the quality of answers obtained by LLaMA3 may be high or low, which makes it beneficial to use these answers to analyze the GPT-4o-as-a-Judge. Second, if GPT-4o is adopted to infer the answers, using GPT-4o to judge again has potential judgment biases. We use the training set of GSM8K as the sampling pool of in-context samples and use the first 200 samples of the test set in GSM8K for experiments in this paper. Specifically, we randomly sample in-context examples for each ${K}$-shot prompt in each test data for reliable results. To ensure that using more shots provides additional information, any ${K}$-shot ICL prompt in our experiments includes all in-context examples from prompts with less than ${K}$ examples (all examples in the training set).
Furthermore, inspired by Counting-Stars Song et al. (2024), even if too many shots are provided, the GPT-4o may not be able to utilize all of them. Because, in the Counting-Stars benchmark, when the number of pieces of evidence reaches 32, long-context LLMs (e.g., GPT-4 Turbo, Gemini 1.5 Pro, and Claude3 Opus) may no longer accurately obtain all of them. Therefore, adding more many-shot in-context examples is probably not captured by LLMs for learning as a reference to judge. However, the many-shot regime and Counting-Stars are substantially different, so there is no noise from "haystack". Hence, we set the maximum number of the in-context examples to 128, i.e., 128-shots.
In the experiments, we use GPT-4o with public API access, and the specific endpoint is “gpt-4o-2024-05-13”. In addition, for the convenience of introduction, when comparing zero-shot and many-shot regimes, we uniformly refer to the few-shot and many-shot regimes as the many-shot regime.
Figure 7: Win rate compared between zero-shot and many-shot regimes. Here, A and A represent the results of the GPT-4o judge in the many-shot regime, and B and B represent the results of the GPT-4o judge in the zero-shot regime. The designed prompts are shown in Table 2. Notice that "[[C]]" denotes a tie.
Figure 8: Comparison evaluation results after mitigating the biases via different approaches.
3.2 Consistency Evaluation
We investigate the consistency between different versions of evaluation results generated by the GPT-4o judge. Here, the single answer grading evaluation results can be used to compare the consistency between different versions. As shown in Figure 3, the bar corresponding to "$2^{n}$" on the x-axis represents the number of samples with consistent and inconsistent ratings in comparing evaluation results obtained twice using GPT-4o-as-a-Judge in the many-shot Reinforced ICL regime. From the results, consistency improves as we increase the number of shots provided as in-context examples during inference.
Recent studies Wang et al. (2023b); Zheng et al. (2023); Chen et al. (2024) have shown that the judgment performance of GPT-4 as a judge is highly in agreement with those of humans. However, both human and LLM judgments are subject to potential biases Chen et al. (2024). By analyzing prior studies, we suppose it unnecessary to obtain utterly accurate evaluation results (because this is difficult) by using LLMs as judges. It is only required to ensure that the evaluation results are highly consistent multiple times so that the single answer grading of GPT-4o-as-a-Judge may be effective. From all results in this work, we find that the many-shot ICL examples help the judgment of LLMs more consistently, which is essential. We consider that the main reason may be that the many-shot in-context examples mitigate the potential biases of the GPT-4o judge in the zero-shot regime.
Meanwhile, we also implemented experiments in the many-shot Unsupervised ICL, but the results show that this regime may be unsuitable in the scenario of acting as a judge because the problem of each sample is the same, but the scored questions and answers are different. From the results in Figure 4, Figure 5, and Figure 6, it can be seen that the consistency corresponds to the number of appended in-context samples with model-generated rationales but nearly not to the number of in-context examples without model-generated rationales.
In addition, both the few-shot and many-shot regimes are sensitive to the selection and order of in-context examples. In the experimental setting of this paper, almost no test data will have the same in-context examples. Even under this condition, the evaluation results also show a high consistency, demonstrating the effectiveness of many-shot ICL in helping GPT-4o-as-a-Judge.
3.3 Quality Evaluation
After the consistency evaluation, an important question arises about: Does a high consistency refer to high-quality judgments? Around this issue, we pair-wise compare model-generated judgment rationales between zero-shot and many-shot regimes using the designed prompts shown in Table 2.
From Figure 7 Compare(A, B), we can find that the evaluations obtained in the many-shot regime are significantly better than those in the zero-shot regime. However, as mentioned before, the positional bias of GPT-4o-as-a-Judge may cause these results, so we performed a second comparison by swapping the positions. As shown in Figure 7 Compare(B, A), we observe that as in-context examples increase, the judgment results in the many-shot regime gradually turn around, that is, a higher winning rate. To obtain fairness results, we integrate the above two results (Compare(A, B) and Compare(B, A)), as shown in Figure 8. It can be seen that after mitigating the positional bias, the judgment quality in the many-shot regime is still better than that in the zero-shot regime.
3.4 Revealing Symbol Bias
Inspired by Song et al. (2024), LLMs revealed a preference for specific symbols under stress testing (details are described in the Appendix). Therefore, an interesting question arises about: Does the GPT-4o prefer to choose the answer with the symbol A rather than B? To verify this conjecture, we swap the answers corresponding to symbols A and B, as shown in Figure 2. From Figure 7 Compare(A, B) and Compare(B, A), it can be seen that the results are different from the above experiments. Actually, the results of Compare(A, B) and Compare(A, B) should be similar, and the results of Compare(B, A) and Compare(B, A) should be similar. This phenomenon shows that symbol bias does exist when adopting GPT-4o-as-a-Judge.
Recent research Wang et al. (2023b) integrates the evaluation results of Compare(A, B) and Compare(B, A) to mitigate the positional bias, which motivates us to incorporate the evaluation results of Compare(A, B) and Compare(B, A) to reduce symbol bias. As presented in Figure 8, it can be seen that as in-context examples increase, the higher the win rate of the many-shot regime, which further verifies the effectiveness of the many-shot regime in helping GPT-4o-as-a-Judge.
Furthermore, there is an interesting hypothesis that the phenomenon of the positional bias we think is not actually caused by the positions of answers, instead of the symbols of answers. In other words, there may be no positional bias in using LLMs as judges in pairwise comparison but rather the symbol bias. The experimental results from Figure 8 Compare(A, B) and Compare(A, B) can demonstrate this hypothesis well.
4 Related Work
LLMs have exhibited remarkable general generation capabilities, positioning themselves as powerful assistants Zhao et al. (2023); OpenAI (2023). With the rapid progression of LLMs, evaluating their proficiency in adhering to human instructions is imperative. Given the advanced capabilities of LLMs, researchers have begun adopting these models to judge the performance of LLMs in following human instructions Koo et al. (2023); Liusie et al. (2023); Liu et al. (2023); Zhu et al. (2023); Lu et al. (2023); Fu et al. (2023b); Zheng et al. (2023); Wang et al. (2023b); Chen et al. (2024). Notably, the evaluation paradigm introduced by Zheng et al. (2023) has gained widespread adoption. However, LLMs as judges are revealed to have potential biases Wang et al. (2023b); Chen et al. (2024), which leads to higher uncertainty and inconsistency during the evaluation using LLMs, questioning the validity of LLM-as-a-Judge.
5 Conclusion
In this work, we mainly study whether many-shot ICL helps long-context LLMs as judges, such as GPT-4o. To this end, we designed several experimental prompts, e.g., many-shot Reinforced and Unsupervised ICL. Experiments show that many-shot ICL can help the GPT-4o judge improve the consistency and quality of evaluation. Meanwhile, we also revealed another bias when LLMs act as judges, symbol bias, and further proposed mitigation approaches. In summary, we preliminarily verified the effectiveness of using many-shot ICL to assist the GPT-4o judge in single answer grading.
6 Limitations
Considering the trade-off between costs and benefits, we do not verify too many in-context examples in the experiments, such as thousands of examples. Combining Figures 3 and Figure 8, it is not difficult to see that when the number of in-context examples increases to 64 and 128, although the consistency no longer improves, the evaluation quality has increased significantly. In addition, we consider that using GPT-4o-as-a-Judge in the many-shot regime is another evolution of the weak-to-strong strategy Burns et al. (2023), which uses many zero-shot judgment results to help GPT-4o generate a better one. As the long-context capabilities of LLMs improve, adding more in-context examples may reveal more valuable phenomena for studying long-context LLMs as judges in the future.
Appendix A Appendix
A.1 Many-Shot Unsupervised ICL
Similar to Agarwal et al. (2024), the default prompt of many-shot Unsupervised ICL is composed of three main components:
  • 1.
    A preamble, such as, “You will be provided questions similar to the ones below:”.
  • 2.
    A list of unsolved inputs or problems.
  • 3.
    A few-shot prompt with outputs for the desired output format.
To require the GPT-4o judge outputs the desired output format, we append 4-shot in-context examples with answers for many-shot Unsupervised ICL. The context length of a single test sample with $K$-shot in-context examples is presented in Table 4. As mentioned before, considering the trade-off between resources and effects, in the experiments of this paper, we do not introduce too many in-context examples, such as thousands of examples.
A.2 Symbol Bias
Previous work Song et al. (2024) has found that when an LLM is stress-tested, it is easy to output some wrong information, which may be an increasing array related to the test data or an English alphabet sequence starting with "A". This phenomenon shows that LLMs favor answers with the symbol "A" rather than the symbol "B" (or the symbol "1" instead of the symbol "2").
A.3 Dataset
GSM8K [^1] https://huggingface.co/datasets/openai/gsm8k is previously introduced by Cobbe et al. (2021), which comprises 8.5K high-quality grade school math problems meticulously crafted by human problem writers. This dataset is divided into 7.5K training and 1K test problems. Specifically, each problem typically requires between 2 to 8 steps to solve, primarily involving a sequence of elementary calculations using basic arithmetic operations (addition, subtraction, multiplication, and division). The problems are designed so that a proficient middle school student can solve each one. Furthermore, the problem-solving task requires models to solve problems with model-generated rationales, which may be challenging to evaluate. We use LLaMA3-70B [^2] https://github.com/meta-llama/llama3 to create model-generated rationales for GSM8K.
Now, I am going to give you a series of demonstrations of Problems and Solutions. When you respond, respond only with the Solution of the final Problem, thinking step by step. Problem Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response. Please rate the response on a scale of 1 to 10 by strictly following this JSON format: {"rating":"", "reason":""}. The "rating" should be as objective as possible. The "reason" denotes a comprehensive explanation of your rating, which should consider factors such as helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response should also be considered. Only return the JSON results and do not give any explanation. Question {question1} Response {response1} Solution {solution1} Problem Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response. Please rate the response on a scale of 1 to 10 by strictly following this JSON format: {"rating":"", "reason":""}. The "rating" should be as objective as possible. The "reason" denotes a comprehensive explanation of your rating, which should consider factors such as helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response should also be considered. Only return the JSON results and do not give any explanation. Question {question2} Response {response2} Solution {solution2} Problem Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response. Please rate the response on a scale of 1 to 10 by strictly following this JSON format: {"rating":"", "reason":""}. The "rating" should be as objective as possible. The "reason" denotes a comprehensive explanation of your rating, which should consider factors such as helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response should also be considered. Only return the JSON results and do not give any explanation. Question {question3} Response {response3} Solution {solution3} Problem Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response. Please rate the response on a scale of 1 to 10 by strictly following this JSON format: {"rating":"", "reason":""}. The "rating" should be as objective as possible. The "reason" denotes a comprehensive explanation of your rating, which should consider factors such as helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response should also be considered. Only return the JSON results and do not give any explanation. Question {question4} Response {response4} Solution {solution4} Problem Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response. Please rate the response on a scale of 1 to 10 by strictly following this JSON format: {"rating":"", "reason":""}. The "rating" should be as objective as possible. The "reason" denotes a comprehensive explanation of your rating, which should consider factors such as helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response should also be considered. Only return the JSON results and do not give any explanation. Question {final_question} Response {final_response} Solution
Table 1: Example Prompt with 4-shot in-context examples used for Reinforced ICL.
Please act as an impartial judge and evaluate the quality of the responses provided by two AI assistants to the user question displayed below. You should choose the assistant that follows the user’s instructions and answers the user’s question better. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of their responses. Begin your evaluation by comparing the two responses and provide a short explanation. Avoid any position biases and ensure that the order in which the responses were presented does not influence your decision. Do not allow the length of the responses to influence your evaluation. Do not favor certain names of the assistants. Be as objective as possible. After providing your explanation, output your final verdict by strictly following this format: "[[A]]" if assistant A is better, "[[B]]" if assistant B is better, and "[[C]]" for a tie.
Problem
Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response. Please rate the response on a scale of 1 to 10 by strictly following this JSON format: {"rating":"", "reason":""}. The "rating" should be as objective as possible. The "reason" denotes a comprehensive explanation of your rating, which should consider factors such as helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response should also be considered. Only return the JSON results and do not give any explanation.
Question
{question}
Response
{response}
The Assistant A’s Answer The Assistant A’s Answer The Assistant B’s Answer The Assistant A’s Answer
{Answer-A} {Answer-B} {Answer-A} {Answer-B}
The Assistant B’s Answer The Assistant B’s Answer The Assistant A’s Answer The Assistant B’s Answer
{Answer-B} {Answer-A} {Answer-B} {Answer-A}
Compare(A, B) Compare(B, A) Compare(A, B) Compare(B, A)
Table 2: The prompts used for pairwise comparison.
Prompt(A). The zero-shot prompt for single answer grading in Zheng et al. (2023).
Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response. Begin your evaluation by providing a short explanation. Be as objective as possible. After providing your explanation, please rate the response on a scale of 1 to 10 by strictly following this format: "[[rating]]", for example: "Rating: [[5]]".
Question
{question}
Response
{response}
Prompt(B). The zero-shot prompt for single answer grading in this paper. Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response. Please rate the response on a scale of 1 to 10 by strictly following this JSON format: {"rating":"", "reason":""}. The "rating" should be as objective as possible. The "reason" denotes a comprehensive explanation of your rating, which should consider factors such as helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response should also be considered. Only return the JSON results and do not give any explanation. Question {question} Response {response}
Table 3: Two versions of zero-shot prompts are used for evaluation in this paper.
$2^{0}$ $2^{1}$ $2^{2}$ $2^{3}$ $2^{4}$ $2^{5}$ $2^{6}$ $2^{7}$
Average Context Length in Tokens of A Single Test Sample
Reinforced ICL 0.7K 1.1K 2.2K 3.6K 6.5K 13.6K 26.3K 50.7K
Unsupervised ICL (Append 4-shots) 2.2K 2.6K 3.1K 4.6K 6.9K 11.7K 20.9K 40.8K
Table 4: Context length in tokens for $K$-shot Reinforced and Unsupervised ICL.
You will be provided Problems similar to the ones below: Problem Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response. Please rate the response on a scale of 1 to 10 by strictly following this JSON format: {"rating":"", "reason":""}. The "rating" should be as objective as possible. The "reason" denotes a comprehensive explanation of your rating, which should consider factors such as helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response should also be considered. Only return the JSON results and do not give any explanation. Question {question1} Response {response1} Now, I am going to give you a series of demonstrations of Problems and Solutions. When you respond, respond only with the Solution of the final Problem, thinking step by step. Problem Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response. Please rate the response on a scale of 1 to 10 by strictly following this JSON format: {"rating":"", "reason":""}. The "rating" should be as objective as possible. The "reason" denotes a comprehensive explanation of your rating, which should consider factors such as helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response should also be considered. Only return the JSON results and do not give any explanation. Question {question1} Response {response1} Solution {solution1} Problem Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response. Please rate the response on a scale of 1 to 10 by strictly following this JSON format: {"rating":"", "reason":""}. The "rating" should be as objective as possible. The "reason" denotes a comprehensive explanation of your rating, which should consider factors such as helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response should also be considered. Only return the JSON results and do not give any explanation. Question {question2} Response {response2} Solution {solution2} Problem Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response. Please rate the response on a scale of 1 to 10 by strictly following this JSON format: {"rating":"", "reason":""}. The "rating" should be as objective as possible. The "reason" denotes a comprehensive explanation of your rating, which should consider factors such as helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response should also be considered. Only return the JSON results and do not give any explanation. Question {question3} Response {response3} Solution {solution3} Problem Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response. Please rate the response on a scale of 1 to 10 by strictly following this JSON format: {"rating":"", "reason":""}. The "rating" should be as objective as possible. The "reason" denotes a comprehensive explanation of your rating, which should consider factors such as helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response should also be considered. Only return the JSON results and do not give any explanation. Question {question4} Response {response4} Solution {solution4} Problem Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response. Please rate the response on a scale of 1 to 10 by strictly following this JSON format: {"rating":"", "reason":""}. The "rating" should be as objective as possible. The "reason" denotes a comprehensive explanation of your rating, which should consider factors such as helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response should also be considered. Only return the JSON results and do not give any explanation. Question {final_question} Response {final_response} Solution
Table 5: Example Prompt with 1-shot in-context examples used for Unsupervised ICL.